5. (2 Points) Let -2 -y F (x2+y2+z2)2/3 (2+y2)2/3 (r2+y2+22)2/3 Find the work done by this force field on an F(t) = (1+3t,
Answer #1

-7 -Y Q= -N -5/3 an=+4ay(ny Rn=42(y S13 No Cusl F(Ry-ai+ C Pe-Rjt(Qr-Py)K abave Valus Py, Qz, P, Ra , Qn4Ry Cusl F 0+oj+0k oV3 fCay 2/3 - y Compasens reuroes 1/2 ea fCx)2) = -2 Cスニュジチズ)+C(つ一4 2/3 compasing withwegot C2(7)=C usng , e becommes 1V/2 +C